OpenStudy (anonymous):

Solve 1n 2 + 1n x = 5. Round to the nearest thousandth, if necessary. (please help on a couple ?'s)

OpenStudy (anonymous):

\[ \ln 2+\ln x = 5 \]Start by taking both sides and raising \(e\) to the power of them: \[ \begin{array}{rcl} e^{\ln2+\ln x} &=& e^5\\ e^{\ln2}e^{\ln x} &=& e^5 \\ 2x &=& e^5 \\ x &=& \frac{e^5}{2} \end{array} \]

OpenStudy (anonymous):

= 74.20657955?